OpenStudy (anonymous):

What is the difference between (For example) (1)² and 1²? Or (1)-² and 1-²?

OpenStudy (unklerhaukus):

\[x^{-1}=\frac 1x\] \[x^{-n}=\frac1{x^n}\]

OpenStudy (amistre64):

a negative exponent represents a reciprocal

OpenStudy (anonymous):

(disregard my last comment)

OpenStudy (anonymous):

\[(x)=x\] Therefore\[1^2=(1)^{2}\] \[1^{-2}=(1)^{-2}\]

OpenStudy (unklerhaukus):

\[(x)^m=x^m\] \[(xy)^m=x^my^m\]

OpenStudy (anonymous):

\[1^{anything}=1\]

OpenStudy (amistre64):

\[a^{(-n)} = \frac 1{a^n}\]

OpenStudy (amistre64):

wow, did we all like post at the same time? or is there something peculiar about the OS?

OpenStudy (unklerhaukus):

jam

OpenStudy (unklerhaukus):

knowledge jam

OpenStudy (anonymous):

Seems so

OpenStudy (unklerhaukus):

@KageWisdom do you still have questions?

OpenStudy (anonymous):

Haha. Thanks all. I sort of understand, but am still a bit confused, maybe I shouldn't have used 1 as my example coefficient. The negative exponents seem to be messing me up.

OpenStudy (anonymous):

So, yes, @UnkleRhaukus, I still have questions. If it turns into a fraction or a reciprical (I still don't understand that) then what's next? How would I solve 8 to the -3.

OpenStudy (unklerhaukus):

\[x^{-n}=\frac{1}{x^n}\] \[7^{-2}=\frac{1}{7^2}=\frac 1{49}\]

OpenStudy (anonymous):

Ok, I think I understand now. I think I was doing where I multiplied 7*-7 instead of doing that. It's too early for algebra! :) Thank you guys so much!

OpenStudy (unklerhaukus):

what did you get for eight to the power of negative three?

OpenStudy (anonymous):

1/8 to the 3rd